Given logb(3x) = 25 and logb(27) = k - 3logb(x) then the value of k is what?
*The 'b' after the log that is NOT in brackets is the base of the log.
*Also we have a formula sheet with all of the Log Laws

Answer by lwsshak3(11628)

Given logb(3x) = 25
logb(27) = k - 3logb(x) then the value of k is what?
..
k=logb(27)+3logb(x)
k=logb(3^3)+3logb(x)
k=3logb(3)+3logb(x)
k=3(logb(3)+logb(x)
k=3logb(3x) (convert to single log)
logb(3x) = 25
k=3*25=75 (ans)